OpenStudy (anonymous):

How do you find the sum of a geometric sequence? (-3,18,-108... if there are 8 terms)

OpenStudy (anonymous):

First find the common ratio

OpenStudy (anonymous):

thats when you deivide the second term by the first, right? so it would be -6?

OpenStudy (anonymous):

divide the second term by the first term to find the common ratio

OpenStudy (anonymous):

Yes,-6 is correct

OpenStudy (anonymous):

okay, then what?

OpenStudy (anonymous):

Now, use this `last term = a (common ratio ^ (number of terms - 1))`

OpenStudy (anonymous):

okay, i got 839808 but thats not an answer i have to choose from.

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

(-3,18,-108... if there are 8 terms) i think now you should do this -108 x -6 = 648 648 x -6 = -3888 -3888 x -6 = 23328 23328 x -6 = -139968

OpenStudy (anonymous):

so now -139968 = -3 times - 6 ^n-1

OpenStudy (anonymous):

what are your options?

OpenStudy (anonymous):

my options are -719835, 719835, 119973, and -119973

OpenStudy (anonymous):

I think its 719835

OpenStudy (anonymous):

Certainly

OpenStudy (anonymous):

Great, thank you